FANDOM


Apollo11EagleNotebookPg065
<Brief page summary to go here.>

Grumman Aircraft Engineering Corporation

LOG

Page No 73115

Engineer

S. SARBELLO

Project

LM-5

Location

ACE 3

Time

NIGHT SHIFT

Title

Date

8/15-16/68

1.RERUN OF SEQ 077-000 TO SEQ 077-020-03 STILL SHOW 1.5 TO 2.0 OHMS BETWEEN X-LUNAR & D - BUSSES AFTER LUT/CSM BOB LSK 560-1120 WAS DEMATED FROM VEHICLE J92.

2. HECHT'S REQUEST TO DEMATE ALL X-LUNAR RELATED CONNECTORS IN PANEL #11 & #16 WAS REJECTED BY JERRY GOLDMACHER. LM-5 VEH. MGR TO GIVE FURTHER DIRECTION.

3. REMATED PTMU P722 TO J101 EMI BOX AND TMA MODE SW "ON". ALSO REMOVED 28 VDC PWR FROM LM/CSM BOB 1119 AND PULL "CSM/PYRO PWR" CB ON 81060.

4.SOLD TDR #32.

5. ECA #4, TERMINALS 1, 2 & 4 HAVE POTTING COMPOUND REMOVED (BY DRY PICKING METHOD)

6. STE HAS SHUT DOWN VGPS 82140 VIA OCP 70010.




Engineer

A. Hecht / E. Homburger

Project

LM 5

Location

Time

Day Shift

Title

Date

8/16/68


(1) Ran TDR #33 steps 22 thru 25 which prove that lack of isolation stems from 1 of 4 equipments presently removed from vehicle: IMU, PTA, RREA & RRAA. Reading was 300 Ω / 250 Ω. Got PTMU Alarm when completed. Called STO who said that alarm was reset & no alarm condition exists.

(2) Informed STE that EPS shall be notified when each of above boxes is installed, so that further checks ca be made then.

(3) 1820: E. Dowse's son Jimmy was notified that Ed is not expected to work on Sunday.

(2A) Ran TDR33, steps 26-32.

Previous page (64) index Next page (66)